# Sensitive 3mm Imaging of Discrete Sources in the Fields of tSZ-Selected Galaxy Clusters

### v1

@@ 3.2 Sources found
An end-to-end test of our pipeline was carried out by injecting fake sources into the raw MUSTANG2 timestreams for each cluster. Sources had random locations and amplitudes between 10 and 30 mJy. Using the same processing steps described above, we were able to recover the amplitudes and locations of all sources without bias, thus confirming the fidelity of our mapmaking. As a null test, our pipeline was run looking for peaks with amplitudes below $-5\sigma$ ($-7\sigma$ in the outer portions of the maps). No negative peaks were found so we can conclude that the vast majority of our candidate sources are real. A further null test was made by searching for sources in the noise maps. For the 243 clusters observed under good conditions, no positive or negative sources were found confirming the quality of this data.
@@
Figure 3: The fluxes and distances from the cluster centers of sources found in our survey of ACT DR5 clusters. Shown in red is the response of the ACT DR5 cluster pipeline to a compact source. Solid parts of this line are regions where a source reduces the measured Compton-$y$ while in the dashed region a source increases the measured value. The black dashed line represents the 50% completeness limit of our survey below which our constraining power on the statistical properties of sources becomes unreliable. As might be expected with a tSZE selected sample there are no bright sources in the center but even so there is a deficit of sources below 10 mJy.
@@
Overall we found only 24 sources (Table 2, Figure 3) a factor of three less than expected by scaling from the ad-hoc sample in Dicker et al. (2021). Some of this discrepancy can be explained by the completeness of our survey at lower flux levels. Although 1 mJy is over 5 times the map noise, some fraction of 1 mJy sources will coincide with negative noise fluctuations and fall below our detection thresholds (Malmquist bias). This effect will also increase the measured flux of a population of sources near a survey limit requiring that the measured fluxes be de-boosted. To better characterize this effect, further simulations were carried out adding sources to the final MIDAS maps. By using a large number of sources with fluxes between 0.5 and 10 mJy it was possible to calculate the completeness and bias in measured flux as a function of radius from the cluster center and measured flux. However, even when these effects are taken into account, there are still far fewer sources found in the M2-ACT survey than expected. This is discussed more in Section 5.
@@
Figure 4: A signal to noise (SNR) plot of one of our sources (top) and the corresponding spectral energy density (SED) plot (below). On the SNR plot, the MUSTANG2 source is shown as the red circle along with its detected SNR value. Sources from other point source catalogs are shown as the same symbols used on the SED plots. Although error bars are plotted on the SED plot in many cases they are not visible. When a point source catalog did not contain a counterpart the approximate detection threshold is shown as a triangle with a one sided error bar. The fitted radio spectral index between MUSTANG2 and FIRST (or the median spectral index from D2021 when no FIRST counterpart was found) is shown as the blue dotted line. The green dashed line shows a typical 40K dust spectrum that goes through the MUSTANG2 point. The SPIRE detections or upper limits are mostly below the dust spectrum curves, indicating that MUSTANG2 sources are mostly dominated by radio emission. The WISE and 2MASS data indicate the presence of hot gas in some sources, but this will have a negligible contribution at 90 GHz. Similar plots for all 24 sources found are shown at the end of the pdf.
@@ 4 Source Counterparts
The list of candidate point sources was matched to point source catalogs. If the angular resolution of the catalog was greater than MUSTANG2, a matching radius of ${}^{\prime\prime}$ was used. If the catalog was from a lower resolution survey the resolution of the survey was used. Spectral energy density plots for each source where made – see Figure 4 for an example.
@@ 4.1 Radio counterparts
Searches for radio counterparts were carried out using catalogs made from the 150 MHz TGSS (Intema et al., 2017), the 1.4 GHz FIRST survey (White et al., 1997), and the 2–4 GHz VLASS survey (Cohen et al., 2007). These surveys have detection limits of approximately 24.5 mJy, 1 mJy, and 0.5 mJy respectively. Using a search radius of ${}^{\prime\prime}$ matches were found for 75% of our detected sources. In two cases two counterparts were found within our search radius. As these were within one MUSTANG2 beam the sum of the total integrated flux was used when plotting spectral energy density (SED) plots and calculating spectral indexes for use in Equation 1.
@@
An effort was made to see if the FIRST and VLASS catalogs could be used to predict which sources could have fluxes at 90 GHz large enough to affect Compton-$y$ measurements. For all sources which appeared in both catalogs and within $4^{\prime}$ of the center one of the clusters in our sample, the flux at 90 GHz was predicted assuming a constant spectral index between FIRST and 90 GHz. As shown in Figure 5, there is little difference in either the predicted 90 GHz flux or the spectral index between those sources that are counterparts to MUSTANG2 and those that are not seen. Of the sources that are seen by MUSTANG2, some counterparts predict lower 90 GHz fluxes (possibly due to dust), while others predict higher fluxes indicating a steepening of the spectral index that is often seen in radio sources at higher frequencies (Tucci et al., 2011). Any source variability could also cause discrepancies between measured and predicted 90 GHz fluxes. However, there are a large number of sources with predicted fluxes well above the M2-ACT detection limit that are not seen – far larger than could be explained by source variability (which would cause a random spread). This is best explained by most sources having a steepening of the spectral index towards high frequencies.
@@
Possible changes in spectral index at 1.4 GHz and at 90 GHz and the fact that not every source had a counterpart strongly argues for high resolution follow up at other tSZ frequencies such as 150 GHz. As complete surveys of every cluster are not needed such observations could be made quickly.
Figure 5: 90 GHz fluxes predicted using all pairs of sources from the FIRST and VLASS catalogs within our clusters. Red crosses represent those not detected by MUSTANG2 while blue dots represent detections. For all detections the measured value is shown as a black point and is often higher or lower by an order of magnitude. This failure to predict the correct flux, especially the large number of predicted fluxes above our detection threshold, demonstrates the need for the higher frequency measurements in this paper.
4.2 Higher frequency counterparts
Submillimeter and infrared continuum emission from galaxies is normally well described by a so-called grey-body, with $25~{}{\rm K}\lesssim T\lesssim 50{\rm~{}K}$ and $\beta\sim 1.5$. Available point source catalogs at these frequencies include SPIRE (600–1200 GHz: Schulz et al., 2018b), WISE (12–89 THz: Schulz et al., 2018a), and 2MASS (138–240 THz: Jarrett et al., 2000). Spectral energy density plots in the extended version of Figure 4 show the majority of the MUSTANG2 detections have WISE counterparts. However, searches carried out at random locations on each map showed that over 25% of these matches could be line of sight coincidences. The 2MASS counterparts seen in 50% of MUSTANG2 sources are more reliable and their amplitudes point towards a hotter dust component which would have an insignificant amount of emission at 90 GHz. Not all the clusters in our ACT sample had data from SPIRE but for those that did have coverage, few SPIRE sources were counterparts to MUSTANG2. However, it was possible to place upper limits on the 600–1200 GHz emission. In most cases these limits were far below what would be expected if the MUSTANG2 sources were dominated by cold dust (Figure 4). This constraint along with those from 2MASS implies the majority of the flux MUSTANG2 sees is radio emission, consistent with the findings from Dicker et al. (2021) and Orlowski-Scherer et al. 2021.
{rotatetable*}
@@ 4.2 Higher frequency counterparts
Table 2: Sources found by our survey. None of the sources were strong enough to enable fitting for extended emission so only peak fluxes are quoted.
@@
M2-SJ0201.8--0212 ACT-CLJ0201.6-0211 02h01m46.8s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 1.38$\pm$0.14 ${}^{\prime}$0.90 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0201.7--0212 ACT-CLJ0201.6-0211 02h01m43.3s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 1.46$\pm$0.16 ${}^{\prime}$0.52 FIRST,VLASS,WISE,2MASS
M2-SJ0204.5$+$0320 ACT-CLJ0204.5+0321 02h04m29.3s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 8.00$\pm$0.38 ${}^{\prime}$ 0.27 FIRST,VLASS,WISE,2MASS
M2-SJ0208.5--0259 ACT-CLJ0208.3-0255 02h08m27.7s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 7.48$\pm$0.59 ${}^{\prime}$0.46
M2-SJ0220.8--0333 ACT-CLJ0220.9-0332 02h20m49.5s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 2.93$\pm$0.27 ${}^{\prime}$0.85 TGSS,FIRST,VLASS,WISE
M2-SJ0223.7$+$0019 ACT-CLJ0223.6+0020 02h23m39.3s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 7.44$\pm$1.05 ${}^{\prime}$0.46 WISE
M2-SJ0231.7--0453 ACT-CLJ0231.7-0453 02h31m41.3s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 2.86$\pm$0.27 ${}^{\prime}$0.46
M2-SJ0233.3$+$0446 ACT-CLJ0233.2+0448 02h33m16.5s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 6.63$\pm$0.68 ${}^{\prime}$0.46 WISE
M2-SJ0248.1--0332 ACT-CLJ0248.0-0331 02h48m03.3s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 0.93$\pm$0.07 ${}^{\prime}$0.91 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0248.4--0334 ACT-CLJ0248.3-0337 02h48m21.9s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 3.73$\pm$0.24 ${}^{\prime}$0.20 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0249.6$+$0212 ACT-CLJ0249.6+0210 02h49m36.9s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 1.36$\pm$0.15 ${}^{\prime}$0.46
M2-SJ0839.7--0153 ACT-CLJ0839.5-0150 08h39m40.3s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 7.14$\pm$0.34 ${}^{\prime}$ 0.28 FIRST,VLASS,WISE
M2-SJ0839.7--0150 ACT-CLJ0839.5-0150 08h39m43.4s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 4.21$\pm$0.38 ${}^{\prime}$ 0.37 FIRST,WISE,2MASS
M2-SJ0901.4$+$0259 ACT-CLJ0901.5+0301 09h01m26.1s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 4.55$\pm$0.37 ${}^{\prime}$0.14 FIRST,VLASS,WISE,2MASS
M2-SJ0905.7$+$0434 ACT-CLJ0905.6+0434 09h05m43.9s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 5.80$\pm$0.23 ${}^{\prime}$0.09 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0918.8$+$0213 ACT-CLJ0918.7+0211 09h18m48.8s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 4.87$\pm$0.21 ${}^{\prime}$0.03 FIRST,VLASS,WISE,2MASS
M2-SJ1118.0--0211 ACT-CLJ1117.9-0211 11h17m59.7s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 1.74$\pm$0.17 ${}^{\prime}$0.09
M2-SJ1148.2--0236 ACT-CLJ1148.0-0234 11h48m14.5s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 1.30$\pm$0.18 ${}^{\prime}$1.10 TGSS,FIRST,VLASS,WISE
M2-SJ1152.4$+$0029 ACT-CLJ1152.2+0031 11h52m22.5s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 3.27$\pm$0.29 ${}^{\prime}$0.19 TGSS,FIRST,VLASS
M2-SJ1204.0$+$0127 ACT-CLJ1203.9+0126 12h04m02.2s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 3.54$\pm$0.20 ${}^{\prime}$0.04 FIRST,VLASS,WISE,2MASS
M2-SJ1210.3$+$0224 ACT-CLJ1210.2+0223 12h10m18.0s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 1.81$\pm$0.29 ${}^{\prime}$0.88 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ1220.3--0241 ACT-CLJ1220.4-0240 12h20m18.5s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 11.67$\pm$0.22 ${}^{\prime}$0.60 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ1224.5$+$0211 ACT-CLJ1224.3+0211 12h24m27.9s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 4.04$\pm$0.34 ${}^{\prime}$1.20 VLASS
M2-SJ1401.0$+$0252 ACT-CLJ1401.0+0252 14h01m00.8s ${}^{\circ}$${}^{\prime}$${}^{\prime\prime}$ 1.15$\pm$0.16 ${}^{\prime}$0.08 FIRST,WISE

### v2

@@ 3.2 Sources found
An end-to-end test of our pipeline was carried out. Fake sources were placed in noiseless maps at random locations within $4^{\prime}$ of the cluster centers. These maps were sampled using the telescope pointing information and the results added to the raw time streams. The fake sources had random amplitudes between 10 and 30 mJy. Using the same processing steps described above, we were able to recover the amplitudes and locations of all sources without bias, thus confirming the fidelity of our mapmaking and source finding in the high-SNR limit. To ensure our detection threshold was not too low, our pipeline was run looking for peaks with amplitudes below $-5\sigma$ ($-7\sigma$ in the outer portions of the maps). No negative peaks were found, so we can conclude that the vast majority of our candidate sources are real. A further null test was made by searching for sources in the noise maps. For the 243 clusters observed under good conditions, no positive or negative sources were found, confirming the quality of these data. Repeating these tests with values of $\sigma$ 0.2 lower resulted in multiple false detections in the noise maps and at negative values of $\sigma$. Using values 0.2 higher would have resulted in missing 3 sources. From this, we conclude that our chosen SNR cuts based on Figure 2 are close to optimal.
@@
Figure 3: The fluxes and distances from the cluster centers of sources found in our survey of ACT DR5 clusters. Shown in red is the response of the ACT DR5 cluster pipeline to a compact source. Solid parts of this line are regions where a source reduces the measured Compton-$y$ while in the dashed region a source increases the measured value. The black dashed line represents the 50% completeness limit of our survey, below which our constraining power on the statistical properties of sources becomes unreliable. As might be expected with a tSZE-selected sample, there are no bright sources in the center, but even so there is a deficit of sources below 10 mJy in the central $1^{\prime}$.
@@
Overall, we found only 24 sources closer to the center of the clusters than $4^{\prime}$ and with a significance greater than our detection thresholds (Table 2, Figure 3), a factor of three less than expected by scaling from the ad hoc sample in Dicker et al. (2021). Some of this discrepancy can be explained by the completeness of our survey at lower flux levels. Although 1 mJy is over 5 times the map noise, some fraction of 1 mJy sources will coincide with negative noise fluctuations and fall below our detection thresholds (completeness or Malmquist bias). This effect will also increase the measured flux of a population of sources near a survey limit, requiring that the measured fluxes be deboosted. To better characterize the completeness and bias, further simulations were carried out. A large number of fake sources with fluxes between 0.5 and 10 mJy were placed at random locations in the final MIDAS maps. The fraction of these that were detected gave the completeness of our survey, while the fitted amplitudes gave the bias in the measured fluxes. Both of these are functions of radius from the cluster center and flux. However, even when these effects are taken into account, there are still fewer sources found in the M2-ACT survey than expected. This is discussed more in Section 5.
@@
Figure 4: An SNR plot of one of our sources (top) and the corresponding SED plot (bottom). On the SNR plot, the MUSTANG2 source is shown as the red circle along with its detected SNR value. Sources from other point-source catalogs are shown as the same symbols used on the SED plots. Although error bars are plotted on the SED plot, in many cases, they are not visible. When a point-source catalog did not contain a counterpart, the approximate detection threshold is shown as a triangle with a one-sided error bar. The fitted radio spectral index between MUSTANG2 and FIRST (or the median spectral index from D2021 when no FIRST counterpart was found) is shown as the blue dotted line. The green dashed line shows a typical 40 K dust spectrum that goes through the MUSTANG2 point. The SPIRE detections or upper limits are mostly below the dust spectrum curves, indicating that MUSTANG2 sources are mostly dominated by radio emission. The WISE and 2MASS data indicate the presence of hot gas in some sources, but this will have a negligible contribution at 90 GHz. Similar plots for all 24 sources found are shown at the end of the pdf.
@@ 4 Source Counterparts
The list of candidate point sources was matched to point-source catalogs. If the angular resolution of the catalog was greater than MUSTANG2, a matching radius of 9′′ was used. If the catalog was from a lower-resolution survey, the resolution of the survey was used. Spectral energy density (SED) plots for each source were made — see Figure 4 for an example.
@@ 4.1 Radio counterparts
Searches for radio counterparts were carried out using catalogs made from the 150 MHz TGSS (Intema et al., 2017), the 1.4 GHz FIRST survey (White et al., 1997), and the 2–4 GHz VLASS survey (Gordon et al., 2021). These surveys have detection limits of approximately 24.5 mJy, 1 mJy, and 0.5 mJy, respectively. Using a search radius of 9′′, matches were found for 75% of our detected sources. In two cases, two counterparts were found within our search radius. As these were within one MUSTANG2 beam, the sum of the total integrated flux was used when plotting spectral energy density (SED) plots and calculating spectral indexes for use in Equation 1.
@@
An effort was made to see if the FIRST and VLASS catalogs could be used to predict which sources could have fluxes at 90 GHz large enough to affect Compton-$y$ measurements. For all sources that appeared in both catalogs and within $4^{\prime}$ of the center of one of the clusters in our sample, the flux at 90 GHz was predicted assuming a constant spectral index between FIRST and 90 GHz. As shown in Figure 5, there is little difference in the distributions of either the predicted 90 GHz flux or the spectral index between those sources that are counterparts to MUSTANG2 and those that are not seen. Of the sources that are seen by MUSTANG2, some counterparts predict lower 90 GHz fluxes (possibly due to dust), while others predict higher fluxes indicating a steepening of the spectral index that is often seen in radio sources at higher frequencies (Tucci et al., 2011). Any source variability could also cause discrepancies between measured and predicted 90 GHz fluxes. However, there are a large number of sources with predicted fluxes well above the M2-ACT detection limit that are not seen – far larger than could be explained by source variability (which would cause a random spread). This is best explained by most sources having a steepening of the spectral index toward high frequencies.
@@
Possible changes in spectral index at 1.4 GHz and at 90 GHz and the fact that not every source had a counterpart strongly argue for high-resolution follow-up at other tSZ frequencies such as 150 GHz. As complete surveys of every cluster are not needed, such observations could be made quickly.
4.2 Higher-frequency counterparts
Submillimeter and infrared continuum emission from galaxies is normally well described by a so-called greybody, with $25~{}{\rm K}\lesssim T\lesssim 50{\rm~{}K}$ and $\beta\sim 1.5$. Available point-source catalogs at these frequencies include SPIRE (600–1200 GHz: Schulz et al., 2018), WISE (12–89 THz: Cutri et al., 2012), and the Two Micron All Sky Survey (2MASS, 138–240 THz: Jarrett et al., 2000). {rotatetable*}
@@ 4.2 Higher-frequency counterparts
Table 2: Sources found by our survey. None of the sources were strong enough to enable fitting for extended emission, so only peak fluxes are quoted. A machine readable version of this table that includes the fluxes of counterparts can be found as supplementary material.
@@
M2-SJ0201.8--0212 ACT-CLJ0201.6-0211 02h01m46.8s --021156′′ 1.38$\pm$0.14 1.00 -0.90 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0201.7--0212 ACT-CLJ0201.6-0211 02h01m43.3s --021145′′ 1.46$\pm$0.16 0.14 -0.52 FIRST,VLASS,WISE,2MASS
M2-SJ0204.5$+$0320 ACT-CLJ0204.5+0321 02h04m29.3s $+$031933′′ 8.00$\pm$0.38 2.44 0.27 FIRST,VLASS,WISE,2MASS
M2-SJ0208.5--0259 ACT-CLJ0208.3-0255 02h08m27.7s --025910′′ 7.48$\pm$0.59 3.31 -0.46
M2-SJ0220.8--0333 ACT-CLJ0220.9-0332 02h20m49.5s --033312′′ 2.93$\pm$0.27 1.96 -0.85 TGSS,FIRST,VLASS,WISE
M2-SJ0223.7$+$0019 ACT-CLJ0223.6+0020 02h23m39.3s $+$001851′′ 7.44$\pm$1.05 2.13 -0.46 WISE
M2-SJ0231.7--0453 ACT-CLJ0231.7-0453 02h31m41.3s --045243′′ 2.86$\pm$0.27 0.89 -0.46
M2-SJ0233.3$+$0446 ACT-CLJ0233.2+0448 02h33m16.5s $+$044546′′ 6.63$\pm$0.68 2.99 -0.46 WISE
M2-SJ0248.1--0332 ACT-CLJ0248.0-0331 02h48m03.3s --033140′′ 0.93$\pm$0.07 0.10 -0.91 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0248.4--0334 ACT-CLJ0248.3-0337 02h48m21.9s --033421′′ 3.73$\pm$0.24 3.10 -0.20 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0249.6$+$0212 ACT-CLJ0249.6+0210 02h49m36.9s $+$021215′′ 1.36$\pm$0.15 2.35 -0.46
M2-SJ0839.7--0153 ACT-CLJ0839.5-0150 08h39m40.3s --015233′′ 7.14$\pm$0.34 3.00 0.28 FIRST,VLASS,WISE
M2-SJ0839.7--0150 ACT-CLJ0839.5-0150 08h39m43.4s --015019′′ 4.21$\pm$0.38 3.08 0.37 FIRST,WISE,2MASS
M2-SJ0901.4$+$0259 ACT-CLJ0901.5+0301 09h01m26.1s $+$025926′′ 4.55$\pm$0.37 2.40 -0.14 FIRST,VLASS,WISE,2MASS
M2-SJ0905.7$+$0434 ACT-CLJ0905.6+0434 09h05m43.9s $+$043342′′ 5.80$\pm$0.23 1.24 -0.09 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ0918.8$+$0213 ACT-CLJ0918.7+0211 09h18m48.8s $+$021326′′ 4.87$\pm$0.21 2.17 -0.03 FIRST,VLASS,WISE,2MASS
M2-SJ1118.0--0211 ACT-CLJ1117.9-0211 11h17m59.7s --021107′′ 1.74$\pm$0.17 0.27 -0.09
M2-SJ1148.2--0236 ACT-CLJ1148.0-0234 11h48m14.5s --023603′′ 1.30$\pm$0.18 2.53 -1.10 TGSS,FIRST,VLASS,WISE
M2-SJ1152.4$+$0029 ACT-CLJ1152.2+0031 11h52m22.5s $+$002857′′ 3.27$\pm$0.29 3.14 -0.19 TGSS,FIRST,VLASS
M2-SJ1204.0$+$0127 ACT-CLJ1203.9+0126 12h04m02.2s $+$012642′′ 3.54$\pm$0.20 1.28 -0.04 FIRST,VLASS,WISE,2MASS
M2-SJ1210.3$+$0224 ACT-CLJ1210.2+0223 12h10m18.0s $+$022340′′ 1.81$\pm$0.29 0.28 -0.88 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ1220.3--0241 ACT-CLJ1220.4-0240 12h20m18.5s --024058′′ 11.67$\pm$0.22 1.61 -0.60 TGSS,FIRST,VLASS,WISE,2MASS
M2-SJ1224.5$+$0211 ACT-CLJ1224.3+0211 12h24m27.9s $+$021103′′ 4.04$\pm$0.34 2.37 -1.20 VLASS
M2-SJ1401.0$+$0252 ACT-CLJ1401.0+0252 14h01m00.8s $+$025141′′ 1.15$\pm$0.16 0.81 -0.08 FIRST,WISE
SED plots in the extended version of Figure 4 show that the majority of the MUSTANG2 detections have WISE counterparts. However, searches carried out at random locations on each map showed that over 25% of these matches could be line-of-sight coincidences. The 2MASS counterparts seen in 50% of the MUSTANG2 sources are more reliable, and their amplitudes point toward a hotter dust component that would have an insignificant amount of emission at 90 GHz. Not all the clusters in our ACT sample had data from SPIRE, but for those that did have coverage, few SPIRE sources were counterparts to MUSTANG2. However, it was possible to place upper limits on the 600–1200 GHz emission. In most cases, these limits were far below what would be expected if the MUSTANG2 sources were dominated by cold dust (Figure 4). This constraint, along with those from 2MASS, implies that the majority of the flux MUSTANG2 sees is radio emission, consistent with the findings from Dicker et al. (2021) and Orlowski-Scherer et al. 2021.
Figure 5: The 90 GHz fluxes predicted using all pairs of sources from the FIRST and VLASS catalogs within our clusters. Red crosses represent those not detected by MUSTANG2 while blue dots represent detections. For all detections, the measured value is shown as a black point and is often higher or lower by an order of magnitude. This failure to predict the correct flux, especially the large number of predicted fluxes above our detection threshold, demonstrates the need for the higher-frequency measurements in this paper.